OpenStudy (anonymous):

Use any method to solve the following system of equation4x-2y = -4 x + y = 2

OpenStudy (anonymous):

x+y=2 x=2-y then u plug it in 4(2-y)-2y=-4 8-4y-2y=-4 8-6y=-4 -6y=-12 y=2 x=0

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

these ppl on here dont write back for like a week

hero (hero):

Did you understand how to solve the system @jazzyo

OpenStudy (anonymous):

I got it! Thanks nick123411!